# Learning Lens Blur Fields

### v1

@@ IV Results
We evaluate our method in simulation and on captured data, and perform quantitative and qualitative comparisons to other non-blind PSF estimation methods. We also describe recovered blur fields for a wide range of lenses and devices, ranging from iPhone and Google Pixel cameras to SLR lenses of different varieties.
@@ IV-A Simulated Evaluation
To evaluate our method in simulation, we use calibration patterns blurred with a known PSF. Specifically, we use the same synthetic noise patterns outlined before and blur them with a spatially-invariant parametric model of lens blur for dual-pixel cameras [10]. The focus is adjusted by varying the radius of the PSFs using the thin lens model [69], with a 30 mm focal length, 40 cm calibration target distance, and linear sampling in diopter space. The maximum and minimum defocus blur radii are set to 24 and 0.5 pixels respectively. The resulting focal stack comprises $N_{I}=21$ images for each synthetic noise pattern. Gaussian read noise with a standard deviation of 0.01 is added to the blurred images. Additionally, we simulate a 16-bit capture with Poisson noise. Overall, our noise model incorporates read noise (Gaussian), shot noise (Poisson), and ADC (quantization) noise.
@@
We train on the paired blur-free and blurry images, holding out every other image in the focal stack for validation. The proposed approach is compared to PSFs reconstructed using the methods of Joshi et al. [31] and Mannan et al. [32] across focus settings. Both baseline methods employ constrained least squares optimization, penalizing deviations between patches in predicted and captured data. Mannan et al. include an additional penalty over filtered versions of the predicted and captured data (e.g., to facilitate reconstructing edges).  Figure 3 compares the ground truth PSFs to those recovered by our method and the two baselines.
@@
Qualitatively, our lens blur field recovers PSFs closer to the ground truth: they have smooth variations with far fewer of the high-frequency artifacts that appear in other methods as shown in  Figure 3. We find that the proposed method also significantly outperforms the recovered PSFs in terms of quantitative accuracy, as evaluated using root-mean-square error (RMSE), indicated for each PSF displayed in  Figure 3.
@@
We also evaluate how well the predicted PSFs estimate the blurry calibration patterns and report the peak signal-to-noise ratio, RMSE, and structural similarity between the predicted and ground truth blurry patterns. To assess performance at withheld focal stack positions, we query our MLP at the corresponding focus setting. For the other methods, we perform linear interpolation between the nearest optimized PSFs. We find that our method outperforms the baselines when evaluating on the validation lens positions (i.e., positions not seen during training or optimized explicitly by the baselines) as shown in Table I. The image reconstruction metric is less sensitive to the high-frequency artifacts in the PSF obseved in Joshi et al.’s method [31], and so their approach performs roughly on par with our method in terms of overfitting to the blurry image. S5 in the Supplement also contains additional details on the aliasing behaviors of Mannan & Langer [32] and Joshi et al. [31].
Comparison Chart Training Lens Positions Validation Lens Positions
Proposed Joshi et al. [31] Mannan & Langer [32] Proposed Joshi et al. [31] Mannan & Langer [32]
Training Patterns PSNR $\uparrow$ $32.109\pm 1.201$ $31.829\pm 1.016$ $15.420\pm 0.381$ $30.462\pm 3.850$ $30.797\pm 3.172$ $15.291\pm 0.513$
RMSE $\downarrow$ $0.022\pm 0.004$ $0.025\pm 0.003$ $0.166\pm 0.010$ $0.031\pm 0.023$ $0.031\pm 0.013$ $0.169\pm 0.013$
SSIM $\uparrow$ $0.929\pm 0.072$ $0.950\pm 0.071$ $0.601\pm 0.021$ $0.924\pm 0.066$ $0.944\pm 0.066$ $0.600\pm 0.024$
Validation Patterns PSNR $\uparrow$ $32.051\pm 1.179$ $31.981\pm 1.277$ $16.753\pm 3.189$ $30.396\pm 3.840$ $28.964\pm 4.493$ $16.400\pm 3.086$
RMSE $\downarrow$ $0.022\pm 0.004$ $0.022\pm 0.003$ $0.139\pm 0.079$ $0.031\pm 0.023$ $0.038\pm 0.029$ $0.145\pm 0.077$
SSIM $\uparrow$ $0.924\pm 0.078$ $0.916\pm 0.103$ $0.461\pm 0.290$ $0.919\pm 0.072$ $0.874\pm 0.100$ $0.452\pm 0.278$
@@
TABLE I: Evaluation of reconstruction accuracy on synthetic noise patterns. The proposed approach quantitatively outperforms non-parametric, optimization-based methods in terms of peak signal-to-noise ratio (PSNR) and root-mean-square error (RMSE). Validation patterns have the same frequencies as the training set. Additional evaluations are included in the supplement.
@@
Finally, we also validate our approach using experimental data, as presented in Table II. Specifically, we blur a ground-truth calibration image with our learned blur fields and compare the output to real-world blurry captures. Metrics are computed on both training and a held-out validation dataset, the latter containing calibration patterns not seen during training. Our method generalizes well to the validation dataset, as shown in Table II.
@@ IV-B Comparison to Prior Methods
The proposed method outperforms prior methods for dense stacks as shown in Table III. While a direct comparison of runtime and memory for a single PSF favors prior work, our method scales better with the number of images—a requirement for accurate calibration. For the example shown in Table III, consider optimizing a 5D PSF for a 12 MP (4032x3024) iPhone 12 Pro; if we optimize blur kernels at 75x100 sensor positions (i.e., assume each blur kernel is constant within a 40 pixel area), 15 focus positions, 73x73 blur kernel samples, and 4 RGGB color channels, then Joshi et al. (2008) would require 34 days for optimization, Mannan and Langer (2016) would require 1945 days, and our approach requires 14 hours. Moreover storing this discretely sampled 5D PSF requires 9.3 GB whereas our MLP requires 19.1 MB.
Prior methods were also observed to be more noise-sensitive, often produce estimations with high-frequency artifacts, and deteriorates for large PSFs (see Figure 3 and Supplement S5).
@@ IV-C Comparing iPhone 12 Pro and 14 Pro Lens Blur Fields
Our method can resolve the differences in optical behavior between the wide-angle cameras on two separate iPhone 12 Pro devices (Supplement S2) and two separate iPhone 14 Pro devices (Figure 6). To demonstrate this, we conduct two types of repeatability experiments.
First, we assesses the variation in estimated blur fields for a single camera with successive captures, where the only difference is the acquisition app’s closure, and the camera’s placement remains constant. This first type of evaluation measures the sensitivity of our method to device-centric parameters, such as focus-position repeatability or repeatability of camera software settings. Second, we evaluate the variation in estimated blur fields after closing the acquisition app, unmounting, and remounting the phone, which assesses repeatability of the entire acquisition procedure.
For both experiments, we use each device to capture 4 separate focal stack datasets and train a separate model for each repeatablity investigation (i.e., a total of 8 datasets per iPhone). Figure 6 shows recovered 5D lens blur fields and 2D PSF slices recovered for each device, averaged across the four trials. We find that the standard deviation of estimated PSF values across trials is significantly smaller than the RMS difference between the mean PSFs of each device. Moreover, we see clear differences between the PSFs in the Supplement, especially near the PSF boundaries.
Our method is sufficiently sensitive to uncover such cross-device differences without specialized hardware. Supplement S2 contains additional details on these repeatability experiments.
@@ IV-E Comparing SLR Lens Blur Fields
Our methodology demonstrates robustness against dismounting and remounting of SLR lenses, indicating negligible optical axis misalignment between lens removals. Repeatability experiments were conducted using a Canon EF 50mm f/1.4 lens for separate trials with and without removing and re-attaching the lenses (see Figure 6). In each experiment, we captured five focal stack datasets and trained individual models for each, resulting in a total of ten datasets. Comparing the blur fields between remount and without remount experiments showed insignificant differences.
@@ IV-F Reconstructing Dual-Pixel Lens Blur Fields
Our proposed method enables reconstructing lens blur fields for left- and right-looking pixels on dual-pixel cameras, such as the Google Pixel 4 (shown in  Figure 1). We observe a complex dependency of the PSF on image plane location and focus setting, with the observed blur patterns departing radically from simpler parametric models for dual-pixel PSFs [10]. Notably, we find that PSF centers of mass do not always follow the expected left-right orientations, and the PSFs are significantly skewed at locations near the periphery of the image sensor.
@@ IV-G Additional SLR Lens Blur Fields
We recover lens blur fields for a variety of Canon SLR lenses of quality and type, shown in Figure 7. Our method is sensitive enough to recover various properties of the PSFs, including the octagonal shape of the 50 mm Canon f/1.4 lens (due to the aperture blades), and the irregular shape of the Canon 24–70 mm f/2.8L zoom lens (evaluated at a 50 mm focal settings), which is most likely due to vignetting. The Canon EF 14 mm f/2.8L and Canon EF 50 mm f/1.2L prime lenses show smooth, round defocus blur with minimal spatial variation.
@@
We demonstrate recovering the 6D lens blur field of an iPhone 12 Pro’s wide lens as shown in Figure 4. The experiment illustrates our method’s adaptability to increased spatial dimensions. We see similar variations in the PSF over the image plane and focus distance dimensions. We can also visualize how the PSF varies over the object distance dimension.
Figure 4: 6D lens blur field calibrated for an iPhone 12 Pro wide lens. We show 5D slices of the 6D function, which vary over focus distance (left) and target distance (right), ranging from 20.5 cm to 41.5 cm. While we observe small changes in the PSF over target distance, the most dramatic changes occur over the focus dimension. This is expected as smartphone cameras have a small apertures, demonstrating that a 5D blur field may be a sufficient blur characterization for certain applications.
@@ IV-I Partial-Field-of-View Blur Field Acquisition
Our method successfully acquires a full-sensor blur field, even in scenarios where the display spans a partial field of view of the sensor.  Figure 5 shows an experiment where one blur field was trained on two sets of captures, with the patterns partially observed in each set.
@@ V-A Lens Blur Field Dataset
To facilitate applications of lens blur fields in computational photography and rendering, we make available a dataset of lens blur fields calibrated across a range of smartphone cameras and SLR lenses and camera bodies. The dataset includes all lens blur fields described thus far (a detailed listing of the camera and lenses included in the dataset is provided in the Supplement S2).
@@ V-B Resolution Chart Experiment
A comparison of captures and estimates rendered with the calibrated lens blur field of the iPhone 12 Pro visualized in  Figure 5 is presented in the Supplement. We capture a focal stack of a resolution chart and the captured images are compared with estimated images, showing overall agreement across the focus range (see Figure 8).
V-C Rendering Scenes with Blur
To facilitate synthetic rendering, we create a Blender [70] postprocessing add-on that allows one to take advantage of our pre-trained lens blur field models as part of their rendering pipeline. We demonstrate this technique on synthetic scenes shown in Supplement S6, rendered with the occlusion-aware model of Ikoma et al. [71] (which itself is agnostic to whether Blender or another engine is used). Different lens blur fields result in a diverse set of bokeh patterns and we show an additional result on captured data in Supplement S6.

### v2

@@ IV Results
We evaluate our method in simulation and on captured data, and perform quantitative and qualitative comparisons to other non-blind PSF estimation methods. We also describe recovered blur fields for a wide range of lenses and devices, ranging from iPhone and Google Pixel cameras to varying SLR lenses.
@@ IV-A Simulated Evaluation
To evaluate our method in simulation, we use calibration patterns blurred with a known PSF. Specifically, we use the synthetic noise patterns [69] shown in Fig. 2 and blur them with a spatially-invariant parametric model of lens blur for dual-pixel cameras [10]. The focus is adjusted by varying the radius of the PSFs using the thin lens model [79], with a 30 mm focal length, 40 cm calibration target distance, and linear sampling in diopter space. The maximum and minimum defocus blur radii are set to 24 and 0.5 pixels respectively. The resulting focal stack comprises $N_{I}=21$ images for each synthetic noise pattern. Gaussian read noise with a standard deviation of 0.01 is added to the blurred images. Additionally, we simulate a 16-bit capture with Poisson noise. Overall, our noise model incorporates read noise (Gaussian), shot noise (Poisson), and ADC (quantization) noise.
@@
We train on the paired blur-free and blurry images, holding out every other image in the focal stack for validation. The proposed approach is compared to PSFs reconstructed using the methods of Joshi et al. [31] and Mannan et al. [32] across focus settings. Both baseline methods employ constrained least squares optimization, penalizing deviations between patches in predicted and captured data. Mannan et al. include an additional penalty over filtered versions of the predicted and captured data (e.g., to facilitate reconstructing edges).  Fig. 3 compares the ground truth PSFs to those recovered by our method and the two baselines.
@@
Qualitatively, our lens blur field recovers PSFs closer to the ground truth: they have smooth variations with far fewer of the high-frequency artifacts that appear in other methods as shown in  Fig. 3. We find that the proposed method also significantly outperforms the recovered PSFs in terms of quantitative accuracy, as evaluated using root-mean-square error (RMSE), indicated for each PSF displayed in  Fig. 3.
@@
We also evaluate how well the predicted PSFs estimate the blurry calibration patterns and report the peak signal-to-noise ratio, RMSE, and structural similarity between the predicted and ground truth blurry patterns. To assess performance at withheld focal stack positions, we query our MLP at the corresponding focus setting. For the other methods, we perform linear interpolation between the nearest optimized PSFs. We find that our method outperforms the baselines when evaluating on the validation lens positions (i.e., positions not seen during training or optimized explicitly by the baselines) as shown in Tab. I. The image reconstruction metric is less sensitive to the high-frequency artifacts in the PSF obseved in Joshi et al.’s method [31], and so their approach performs roughly on par with our method in terms of overfitting to the blurry image. Supplement Sec. S-IV contains additional comparisons with neural deconvolution baselines. Results in Tab. I also show that directly obtaining PSF predictions at validation lens positions from the MLP (using the MLP’s interpolation) outperforms a baseline that obtains PSFs predictions only at training lens positions and then uses linear interpolation to estimate the validation lens positions.
Training Lens Positions Validation Lens Positions
Proposed Joshi et al. [31] Mannan & Langer [32] Proposed Joshi et al. [31] Mannan & Langer [32] Lin. Interpolation + Proposed
Training Patterns PSNR $\uparrow$ $32.109\pm 1.201$ $31.829\pm 1.016$ $15.420\pm 0.381$ $30.462\pm 3.850$ $30.797\pm 3.172$ $15.291\pm 0.513$ $29.394\pm 3.607$
SSIM $\uparrow$ $0.929\pm 0.072$ $0.950\pm 0.071$ $0.601\pm 0.021$ $0.924\pm 0.066$ $0.944\pm 0.066$ $0.600\pm 0.024$ $0.898\pm 0.077$
RMSE $\downarrow$ $0.022\pm 0.004$ $0.025\pm 0.003$ $0.166\pm 0.010$ $0.031\pm 0.023$ $0.031\pm 0.013$ $0.169\pm 0.013$ $0.034\pm 0.022$
Validation Patterns PSNR $\uparrow$ $32.051\pm 1.179$ $31.981\pm 1.277$ $16.753\pm 3.189$ $30.396\pm 3.840$ $28.964\pm 4.493$ $16.400\pm 3.086$ $29.351\pm 3.651$
SSIM $\uparrow$ $0.924\pm 0.078$ $0.916\pm 0.103$ $0.461\pm 0.290$ $0.919\pm 0.072$ $0.874\pm 0.100$ $0.452\pm 0.278$ $0.887\pm 0.091$
RMSE $\downarrow$ $0.022\pm 0.004$ $0.022\pm 0.003$ $0.139\pm 0.079$ $0.031\pm 0.023$ $0.038\pm 0.029$ $0.145\pm 0.077$ $0.034\pm 0.023$
@@
TABLE I: Evaluation of reconstruction accuracy on synthetic noise patterns. The proposed approach quantitatively outperforms non-parametric, optimization-based methods in terms of peak signal-to-noise ratio (PSNR) and root-mean-square error (RMSE). Validation patterns have the same frequencies as the training set. MLP interpolation outperforms linear interpolation (denoted as “Lin. Interpolation + Proposed”), where PSFs at validation lens positions are obtained by linearly interpolating the MLP predictions at the training lens positions. The difference is most pronounced near the in-focus position because uniform diopter sampling yields non-uniform metric distance sampling, and PSF size changes become more significant due to a derivative sign change near focus. MLP interpolation effectively handles the resulting complex, nonlinear defocus blur radius model derived from the thin lens equation. Additional evaluations are included in Supplement Sec. S-IV.
@@
Finally, we also validate our approach using experimental data, as presented in Tab. II. We blur a ground-truth calibration image with our learned blur fields and compare the output to real-world blurry captures. Metrics are computed on both training and a held-out validation dataset, the latter containing calibration patterns not seen during training. Our method also generalizes well to the validation dataset.
@@ IV-B Comparison to Prior Methods
The proposed method outperforms prior methods for dense stacks as shown in Tab. III. While a direct comparison of runtime and memory for a single PSF favors prior work, our method scales better with the number of images—a requirement for accurate calibration. For the example shown in Tab. III, consider optimizing a 5D PSF for a 12 MP (4032x3024) iPhone 12 Pro; if we optimize blur kernels at 75x100 sensor positions (i.e., assume each blur kernel is constant within a 40 pixel area), 15 focus positions, 73x73 blur kernel samples, and 4 RGGB color channels, then Joshi et al. (2008) would require 34 days for optimization, Mannan and Langer (2016) would require 1945 days, and our approach requires 14 hours. Moreover storing this discretely sampled 5D PSF requires 9.3 GB whereas our MLP requires 19.1 MB. Prior methods were also observed to be more noise-sensitive, often produce estimations with high-frequency artifacts, and deteriorates for large PSFs (see Fig. 3 and Supplement Sec. S-IV).
@@ IV-C Comparing iPhone 12 Pro and 14 Pro Lens Blur Fields
Our method can resolve the differences in optical behavior between the wide-angle cameras on two separate iPhone 12 Pro devices (Supplement Sec. S-II) and two separate iPhone 14 Pro devices (Fig. 6). To demonstrate this, we conduct two types of repeatability experiments.
First, we assesses the variation in estimated blur fields for a single camera with successive captures, where the only difference is the closing and re-launching the acquisition app, while camera placement remains constant. This first type of evaluation measures the sensitivity of our method to device-centric parameters, such as focus-position repeatability or repeatability of camera software settings. Second, we evaluate the variation in estimated blur fields after closing the acquisition app, unmounting, and remounting the phone, which assesses repeatability of the entire acquisition procedure.
For both experiments, we use each device to capture 4 separate focal stack datasets and train a separate model for each repeatablity investigation (i.e., a total of 8 datasets per iPhone). Fig. 6 shows recovered 5D lens blur fields and 2D PSF slices recovered for each device, averaged across the four trials. We find that the standard deviation of estimated PSF values across trials is significantly smaller than the RMS difference between the mean PSFs of each device. Moreover, we see clear differences between the PSFs in the Supplement, especially near the PSF boundaries. Our method is sufficiently sensitive to uncover such cross-device differences without specialized hardware. Supplement Sec. S-II contains additional details on these repeatability experiments.
@@ IV-E Comparing SLR Lens Blur Fields
Our methodology demonstrates robustness against dismounting and remounting of SLR lenses, indicating negligible optical axis misalignment between lens removals. Repeatability experiments were conducted using a Canon EF 50mm f/1.4 lens for separate trials with and without removing and re-attaching the lenses (see Fig. 6). In each experiment, we captured five focal stack datasets and trained individual models for each, resulting in a total of ten datasets. Comparing the blur fields between remount and without remount experiments showed insignificant differences.
@@ IV-F Reconstructing Dual-Pixel Lens Blur Fields
Our proposed method enables reconstructing lens blur fields for left- and right-looking pixels on dual-pixel cameras, such as the Google Pixel 4 (shown in  Fig. 1). We observe a complex dependency of the PSF on image plane location and focus setting, with the observed blur patterns departing radically from simpler parametric models for dual-pixel PSFs [10]. Notably, we find that PSF centers of mass do not always follow the expected left-right orientations, and PSFs are significantly skewed at locations near the periphery of the image sensor.
@@ IV-G Additional SLR Lens Blur Fields
We recover lens blur fields for a variety of Canon SLR lenses of quality and type, shown in Fig. 7. Our method is sensitive enough to recover various properties of the PSFs, including the octagonal shape of the 50 mm Canon f/1.4 lens (due to the aperture blades), and the irregular shape of the Canon 24–70 mm f/2.8L zoom lens (evaluated at 50 mm focal setting), which is most likely due to vignetting. The Canon EF 14 mm f/2.8L and Canon EF 50 mm f/1.2L prime lenses show smooth, round defocus blur with minimal spatial variation.
@@
A 6D blur field extends a 5D blur field by including an additional dimension, $d$, for target distance. We demonstrate recovery of the 6D lens blur field for an iPhone 12 Pro’s wide lens in Fig. 4, illustrating our method’s adaptability to increased spatial dimensions. We see less variation in the PSF over target distance dimension than in other dimensions, such as focus position. This behavior is consistent with that of a thin lens, where variations in target distance can be largely predicted from a 5D blur field.
Figure 4: 6D lens blur field calibrated for an iPhone 12 Pro wide lens and the capture setup for 6D blur fields. We show 5D slices of the 6D function, which vary over focus distance (top left) and target distance (top right), ranging from 20.5 cm to 41.5 cm. While we observe small changes in the PSF over target distance, the most dramatic changes occur over the focus dimension. This is expected as smartphone cameras have a small apertures, demonstrating that a 5D blur field may be a sufficient blur characterization for certain applications. The 6D capture setup comprises a monitor, a phone stand, and a translating stage that controls positioning along the target distance dimension.
@@ IV-I Partial-Field-of-View Blur Field Acquisition
Our method successfully acquires a full-sensor blur field, even in scenarios where the display spans a partial field of view of the sensor.  Fig. 5 shows an experiment where one blur field was trained on two sets of captures, with the patterns partially observed in each set.
@@ V-A Lens Blur Field Dataset
To facilitate applications of lens blur fields in computational photography and rendering, we make available a dataset of lens blur fields calibrated across a range of smartphone cameras and SLR lenses. The dataset includes all lens blur fields described thus far (a complete listing of cameras and lenses included in the dataset is provided in Supplement Tab. S-III).
@@ V-B Resolution Chart Experiment
A comparison of captures and estimates rendered with the calibrated lens blur field of the iPhone 12 Pro visualized in  Fig. 5 is presented in the Supplement. We capture a focal stack of a resolution chart and the captured images are compared with image restorations, showing overall agreement across the focus range (see Fig. 8).
V-C Improved Image Restoration with Blur Fields Compared to Seidel Model
We compare the proposed lens blur fields with the parametric Seidel model by evaluating their performance for image restoration on a real-world capture of a Siemens Star test pattern using an iPhone 12 Pro wide camera (Tab. II). After constructing the lens blur field, we follow Kohli et al. [80] to fit a Seidel model—with spherical, coma, astigmatism, field curvature, and defocus coefficients—to the PSFs recovered by the blur field by randomly sampling 32 image plane locations per lens position. The Seidel coefficients are optimized via gradient descent to match the lens blur field PSFs at the sampled locations using 300 iterations with a learning rate of $0.01$, the default parameters in Kohli et al. [80]. For image restoration, we use the same optimization approach as for PSF estimation (Sec. II-B). We solve for a restored image that, when convolved with the respective blurs, produces a synthetic blurry image. We then compute a loss between this synthetic image and the real-life blurry image.
We compare the resulting restorations and find that blur field PSFs yield superior results. Qualitative results in Fig. 8 demonstrate that the lens blur field better preserves high-frequency details in the Siemens Star pattern, reconstructing longer portions of the star’s arms than do restorations using PSFs from the Seidel model. A direct comparison with the ground truth Siemens Star in Tab. IV shows that restorations using blur field PSFs are superior. Furthermore, the synthetic blurry image generated by convolving the restored Siemens Star with blur field PSFs more closely matches the real-life image than the synthetic image generated using Seidel PSFs.
Restoration Results Siemens Star Restoration Synthetic Blurry Image
Red Green 1 Green 2 Blue
Proposed PSNR $\uparrow$ 8.900 20.335 20.392 21.479 17.700
SSIM $\uparrow$ 0.293 0.747 0.780 0.786 0.600
RMSE $\downarrow$ 0.353 0.050 0.054 0.047 0.059
Seidel Coefficients PSNR $\uparrow$ 7.415 20.184 19.660 20.528 17.754
SSIM $\uparrow$ 0.127 0.534 0.588 0.595 0.515
RMSE $\downarrow$ 0.421 0.051 0.059 0.053 0.058
TABLE IV: Restoration results for both the estimated Siemens Star and the synthetic blurry image obtained by convolving the estimated Star with either the blur fields PSF or the Seidel PSF. Both the restored Siemens Star and the synthetic blurry image show that the blur field PSFs recover real-life effects more effectively than the Seidel model.
V-D Rendering Scenes with Blur
To facilitate synthetic rendering, we create a Blender [81] postprocessing add-on that allows one to take advantage of our pre-trained lens blur field models as part of their rendering pipeline. We demonstrate this technique on synthetic scenes shown in Supplement Sec. S-V, rendered with the occlusion-aware model of Ikoma et al. [82] (which itself is agnostic to whether Blender or another engine is used). Different lens blur fields result in a diverse set of bokeh patterns and we show an additional result on captured data in Supplement Sec. S-V.